OpenStudy (icalibear):

Write an equation in standard form for the points (5,1) and (-3,4)?

OpenStudy (apoorvk):

For a line passing through two points say, A\((x_1,y_1)\) and B\((x_2,y_2)\), the equation is: \[\frac{y-y_1}{x-x_1} = \frac{y_1 - y_2}{x_1 - x_2}\] Can you use this for the points given to you?

OpenStudy (icalibear):

\[y -1/ x - 5 = 1 - 4/ 5 - -3 \]

OpenStudy (icalibear):

idk how to do the division bar on there

OpenStudy (apoorvk):

no worries, you can just use brackets. And yes, that's correct! just simplify it a bit!

OpenStudy (icalibear):

how do i simplify it 0.o

OpenStudy (apoorvk):

Hmm. \[\frac{y−1}{x−5}=\frac{1−4}{5−(−3)}\] okay, so what's '1-4' and what's '5-(-3)'?